OpenStudy (calculusxy):

In a certain coffee shop, a large latte costs 2 dollars more than a medium latte, and a medium latte costs 2 dollars more than a small latte. If 12 small lattes cost 2x dollars, how much will 5 large and 7 medium lattes cost? A. 2x + 24 B. 2x + 34 C. 2x + 48 D. 12x + 24 E. 24x + 3

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what does x represent?

OpenStudy (calculusxy):

I don't know. That is confusing me.

jimthompson5910 (jim_thompson5910):

cost of large latte? medium? or small?

OpenStudy (calculusxy):

Can we assume that x=6?

jimthompson5910 (jim_thompson5910):

where are you getting that?

OpenStudy (calculusxy):

nvm

jimthompson5910 (jim_thompson5910):

S = cost of small M = cost of medium L = cost of large

jimthompson5910 (jim_thompson5910):

`a large latte costs 2 dollars more than a medium latte` so L = M+2 agreed?

OpenStudy (calculusxy):

yes

jimthompson5910 (jim_thompson5910):

` medium latte costs 2 dollars more than a small latte` using the variables I set up, what equation do you get?

OpenStudy (calculusxy):

S = x M = x + 2 L = 2 + (x + 2) or 4 + x

jimthompson5910 (jim_thompson5910):

S = unknown (not x) M = S + 2 L = M + 2 = (S+2) + 2 = S + 4 `12 small lattes cost 2x dollars` 12S = 2x S = (2x)/(12) S = x/6 `how much will 5 large and 7 medium lattes cost?` 5L + 7M = 5(S+4) + 7(S+2)

jimthompson5910 (jim_thompson5910):

the next step is to plug S = x/6 into 5(S+4) + 7(S+2)

OpenStudy (calculusxy):

so we are learning about plugging in numbers. can we work around that level?

jimthompson5910 (jim_thompson5910):

you mean replacing S with some number like 7?

OpenStudy (calculusxy):

yeah

jimthompson5910 (jim_thompson5910):

how would you get to the answer? each answer has an x in it

OpenStudy (calculusxy):

so can we like plug in 7 for x?

OpenStudy (calculusxy):

I am really confused on this.

jimthompson5910 (jim_thompson5910):

Do you see how I got 5(S+4) + 7(S+2) ?

OpenStudy (calculusxy):

yes

jimthompson5910 (jim_thompson5910):

\[\Large 5(S+4) + 7(S+2) = 5(\frac{x}{6}+4) + 7(\frac{x}{6}+2) \] \[\Large 5(S+4) + 7(S+2) = 5(\frac{x}{6})+5(4) + 7(\frac{x}{6})+7(2) \] \[\Large 5(S+4) + 7(S+2) = \frac{5x}{6}+20 + \frac{7x}{6}+14 \] I'll let you simplify

OpenStudy (calculusxy):

12x/6 + 34

jimthompson5910 (jim_thompson5910):

12x/6 simplifies to ???

OpenStudy (calculusxy):

2x

OpenStudy (calculusxy):

THANK YOU!!!

jimthompson5910 (jim_thompson5910):

so 2x+34 is the answer

jimthompson5910 (jim_thompson5910):

np

OpenStudy (calculusxy):

I have another question

OpenStudy (calculusxy):

If \[a = 2x + 3 \] and \[b = 4d\] what is b in terms of a? A. \[(a - 3)^2\] B.\[2(a - 3)^2\] C. \[\frac{ (a-3)^2 }{ 4 }\] D. \[\frac{ (a + 3)^2 }{ 4 }\] E. \[a + 3\]

jimthompson5910 (jim_thompson5910):

b = 4d? is that a typo?

OpenStudy (calculusxy):

Yes sorry \[b = 4d^2\]

jimthompson5910 (jim_thompson5910):

and a = 2x + 3 is correct? or no?

OpenStudy (calculusxy):

that's correct

jimthompson5910 (jim_thompson5910):

I don't see how to connect a and b. They have no variables in common

OpenStudy (calculusxy):

i have checked with many numbers none of which seems to work

jimthompson5910 (jim_thompson5910):

something seems missing

OpenStudy (calculusxy):

i will check on that with my teacher. thank you :)

jimthompson5910 (jim_thompson5910):

np